OpenStudy (anonymous):

Solve 4x + 10y = 94 y = 3x – 28

OpenStudy (anonymous):

system of equations!

OpenStudy (anonymous):

Try substituting the second equation to the first one.

OpenStudy (anonymous):

You have second equation as : \(y = 3x - 28\) So, put this value in first equation: \(4x + 10y = 94\) Put that value of y in this equation you will get x...

OpenStudy (theviper):

4x+10(3x-28)=94 =>4x+30x-280=94 =>34x=94+280 =>34x=374 =>x=374/34=???

OpenStudy (theviper):

Put the value of 'y' in eq 1 & get 'x' then find 'y'

OpenStudy (anonymous):

\[x = 11 \] \[y = 5\]